OpenStudy (anonymous):

I have this question that I cannot solve so if you can please be kind and tell me how to do it Jose paid $45 to rent a canoe. The cost for renting a canoe was $12 per hour and Jose used a coupon for 15% off the total cost. How many hours did he rent the canoe for?

OpenStudy (anonymous):

$45= 85% of the total price that Jose paid

OpenStudy (anonymous):

So you divide $45 by 85 to find out 1%. That= \[\frac{ 9 }{ 17 }\]

OpenStudy (mathmale):

If Jose got 15% off, then he paid 85% of the usual price. Figure out what the usual price would be: \[\frac{ $12 }{ hour }*x(hours)\]

OpenStudy (anonymous):

Ok thank you. I think I got it now! :D

OpenStudy (anonymous):

so you times \[\frac{ 9 }{ 17 } by 100 \to find out what he would of payed without discount\]

OpenStudy (mathmale):

With that 15% discount, Jose paid $45:\[0.85(12x) = $45\]Solve this equation for x.